# Quantifying interdisciplinary synergy in higher STEM education

### v2

@@ 4.4 The CSS network analysis
The CSS between discipline pairs in Eq. (13) forms a weighted matrix or network representing the level of educational synergy between disciplines. As the number of disciplines is $51$, even a visual inspection compared to the hierarchical structure of disciplines provided by the data set [15] is plausible. We show the $51\times 51$ heatmap of the CSS for all of the STEM discipline pairs in Fig. 6. Overall, large CSS values are prominent among closely related engineering disciplines, hierarchically close to each other according to the standard classification in higher levels (forming dark block diagonal patterns) [15], e.g., the chemical engineering cluster, the computer science cluster, electrical engineering cluster, the mechanical engineering cluster, and the material engineering cluster. In contrast, natural science disciplines do not show large CSS values to each other (the lack of dark block diagonals in the natural science region), and they have large CSS values with discipline-specific engineering counterparts, e.g., (life science $\otimes$ biotechnology) and (physics $\otimes$ semiconductor engineering). This connection pattern is an archetypal mesoscale structural example of core-periphery structures [33], with multiple cores composed of separate modules [34].
Figure 6: The heat map of CSS values between the $51$ disciplines (the lowest-level classification) in the data set. The rows and columns are sorted by the intermediate-level classification shown in the upper hierarchical level drawn on the left [15].
Figure 7: The community (a) and core-periphery (b) structures in the CSS network, where the size of a node (discipline) is proportional to the number of departments in the discipline. (a) The fraction of colors represents the community membership [35] of each node, from $1000$ community ensembles with the resolution parameter $\gamma=1$. (b) Darker colors represent larger core score [33] values for nodes.
@@ 5 Discussion and conclusion
This research was supported by the National Research Foundation of Korea (NRF) under the grants RS-2021-NR061247 (GG and SHL) and RS-2022-NR069864 (JY).
@@ Appendix C Community detection in appropriate resolution
Figure 8: The PaI [26] values (left) and the average number of communities (right) as the functions of the resolution parameter $\gamma$ of the modularity function, as a result of $100$ independent runs of the Louvain algorithm [25] applied to the DSN in Sec. 3 for each $\gamma$ value. The points are the mean values, with the shades around them representing the standard deviation.
The resolution parameter $\gamma$ in the modularity function [37, 38] controls the overall scale of communities, where the modularity-maximization process with larger values of $\gamma$ tends to detect smaller communities. Although this flexibility provides a nice knob-like adjustable parameter for the scale of detected communities, it is nontrivial to decide which resolution yields the most meaningful or statistically reliable communities. In Ref. [26], the authors suggest a method to extract the range of $\gamma$ corresponding to the most self-consistent detection results, namely, the regions of locally minimum values of the partition inconsistency (PaI). We apply this method to our DSN with the Louvain algorithm [25], and as shown in Fig. 8, we find there values of $\gamma=0.56$, $1.12$, and $1.77$, corresponding to the locally consistent communities, so we take these values in addition to the trivial case $\gamma=0$ in Sec. 3 and Fig. 3.
@@ 6 Semiconductor Testing and Reliability 3
Table S2: We show the curriculum design process, beginning with the seed courses at the top and adding others with the criterion described in the main text (listed in the order of addition). The courses in boldface are common in all three curricula, highlighting their central role in integrating chemistry and polymer engineering disciplines. The CSS value for this combination is $0.03499$.
@@ 2 Polymer Physics and Chemistry 3 Polymer Physics and Chemistry 3
Table S3: We show the curriculum design process, beginning with the seed courses at the top and adding others with the criterion described in the main text (listed in the order of addition). The CSS value for this combination is 00.

### v3

@@ 4.4 The CSS network analysis
The CSS between discipline pairs in Eq. (13) forms a weighted matrix or network representing the level of educational synergy between disciplines. As the number of disciplines is 515151, even a visual inspection compared to the hierarchical structure of disciplines provided by the data set [15] is plausible. We show the $\times 51$ 51 heatmap of the CSS for all of the STEM discipline pairs in Fig. 6. Overall, large CSS values are prominent among closely related engineering disciplines, hierarchically close to each other according to the standard classification in higher levels (forming dark block diagonal patterns) [15], e.g., the chemical engineering cluster, the computer science cluster, electrical engineering cluster, the mechanical engineering cluster, and the material engineering cluster. In contrast, natural science disciplines do not show large CSS values to each other (the lack of dark block diagonals in the natural science region), and they have large CSS values with discipline-specific engineering counterparts, e.g., (life science $\otimes$ biotechnology) and (physics $\otimes$ semiconductor engineering). This connection pattern is an archetypal mesoscale structural example of core-periphery structures [33], with multiple cores composed of separate modules [34].
Figure 6: The heat map of CSS values between the 515151 disciplines (the lowest-level classification) in the data set. The rows and columns are sorted by the intermediate-level classification shown in the upper hierarchical level drawn on the left [15].
Figure 7: The community (a) and core-periphery (b) structures in the CSS network, where the size of a node (discipline) is proportional to the number of departments in the discipline. (a) The fraction of colors represents the community membership [35] of each node, from 100010001000 community ensembles with the resolution parameter $\gamma=1$. (b) Darker colors represent larger core score [33] values for nodes.
@@ 5 Discussion and conclusion
This research was supported by the National Research Foundation of Korea (NRF) under the grants RS-2021-NR061247 (GG and SHL) and RS-2022-NR069864 (JY), and results of a study on the “Gyeongsangnam-do Regional Innovation System & Education (RISE)” Project, supported by the Ministry of Education and Gyeongsangnam-do.
@@ Appendix C Community detection in appropriate resolution
Figure 8: The PaI [26] values (left) and the average number of communities (right) as the functions of the resolution parameter $\gamma$ of the modularity function, as a result of 100100100 independent runs of the Louvain algorithm [25] applied to the DSN in Sect. 3 for each $\gamma$ value. The points are the mean values, with the shades around them representing the standard deviation.
The resolution parameter $\gamma$ in the modularity function [37, 38] controls the overall scale of communities, where the modularity-maximization process with larger values of $\gamma$ tends to detect smaller communities. Although this flexibility provides a nice knob-like adjustable parameter for the scale of detected communities, it is nontrivial to decide which resolution yields the most meaningful or statistically reliable communities. In Ref. [26], the authors suggest a method to extract the range of $\gamma$ corresponding to the most self-consistent detection results, namely, the regions of locally minimum values of the partition inconsistency (PaI). We apply this method to our DSN with the Louvain algorithm [25], and as shown in Fig. 8, we find there values of $\gamma=0.56$, 1.121.121.12, and 1.771.771.77, corresponding to the locally consistent communities, so we take these values in addition to the trivial case $\gamma=0$ in Sect. 3 and Fig. 3.
@@ 6 Semiconductor Testing and Reliability 3
Table S2: We show the curriculum design process, beginning with the seed courses at the top and adding others with the criterion described in the main text (listed in the order of addition). The courses in boldface are common in all three curricula, highlighting their central role in integrating chemistry and polymer engineering disciplines. The CSS value for this combination is 0.034990.034990.03499.
@@ 2 Polymer Physics and Chemistry 3 Polymer Physics and Chemistry 3
Table S3: We show the curriculum design process, beginning with the seed courses at the top and adding others with the criterion described in the main text (listed in the order of addition). The CSS value for this combination is 0.